OpenStudy (anonymous):

What is the x-coordinate of the vertex of y = |4x – 12| – 3 ?

OpenStudy (amistre64):

divide out that 4 and you should end up with it

OpenStudy (amistre64):

another way is to find the value of x that makes that inner part go to zero

OpenStudy (anonymous):

This is what I did but my answer should only be one 4x-12=0 x=3 that is for x y=0-3 y=-3 Those are the last two of all I did

OpenStudy (amistre64):

3 is correct

OpenStudy (anonymous):

oh so I dont have to fing y using x?

OpenStudy (amistre64):

its only asking for the x part of the vertex

OpenStudy (anonymous):

oh ok thank you!

OpenStudy (amistre64):

youre welcome